# xLSTMAD: A Powerful xLSTM-based Method for Anomaly Detection

### v1

@@ Abstract
Index Terms:
xLSTM, anomaly detection, deep learning, time series
@@ I Introduction
Anomaly detection in multivariate time series is a vital task across domains such as industrial monitoring, cybersecurity, human activity recognition, and space system diagnostics. While statistical techniques and classical machine learning models remain strong baselines, deep learning methods have emerged as powerful alternatives due to their ability to model complex temporal dependencies and high-dimensional patterns. Recent benchmarks like TSB-AD [1] reveal that even simple neural architectures (e.g., CNNs [2], LSTMs [3]) can outperform more complex Transformer variants [4, 5] in anomaly detection tasks—particularly in multivariate settings where expressive temporal modeling is key.
Most state-of-the-art deep learning approaches fall into two broad categories: reconstruction-based models (e.g., autoencoders, VAEs) that detect anomalies through high reconstruction error, and prediction-based models (e.g., LSTMs) that forecast future behavior and identify deviations. However, these models often struggle to balance long-term memory, fine-grained temporal resolution, and computational efficiency [1].
The recently proposed xLSTM architecture offers a promising research direction [6]. As a generalization of LSTM, it combines depth-wise convolutions, residual connections, and multi-scale gating mechanisms in a stackable, efficient structure. xLSTM has demonstrated success in time series forecasting [7], lossless compression [8], and even large-scale language modeling [9], where its linear memory footprint and fast inference make it a viable Transformer alternative. Despite its growing popularity, no prior work has explored xLSTM for anomaly detection.
@@
To rigorously evaluate our approach, we leverage the TSB-AD-M benchmark [1], a comprehensive suite of 17 real-world datasets ranging from industrial control systems (e.g., SWaT, SMD) to spacecraft telemetry (e.g., SMAP, MSL) and human activity monitoring (e.g., OPP), designed to create extremely challenging conditions that are useful to assess the effectiveness of diverse models in diverse conditions. Our results show that the xLSTM encoder-decoder architecture achieves superior performance compared to classical and modern baselines.
@@ II-A Time series anomaly detection
Anomaly detection in time series encompasses a range of classical and modern methods. Traditional statistical approaches assume an underlying generative model of normal behavior and use statistical tests or thresholds to flag outliers. For example, one may fit an ARIMA or seasonal model and then apply hypothesis tests on residuals [10].  [10, 11].
In the deep learning era, reconstruction-based models have become popular for time-series anomaly detection. Autoencoders (AEs) are neural networks trained to reconstruct input sequences, and anomalies are identified by high reconstruction error. Sakurada and Yairi [12] showed that even simple AEs can outperform linear PCA for detecting anomalies in multivariate time series. Modern variants (e.g., denoising AEs, convolutional AEs) follow the same principle.
Generative deep models explicitly learn the probability distribution of normal time-series data. Variational autoencoders (VAEs) and GANs are common in this class. For example, OmniAnomaly [13] uses a recurrent VAE with planar normalizing flows to model multivariate time series.
Recently, transformer and attention-based architectures have been applied to anomaly detection. Transformers process the entire sequence via self-attention, capturing long-range dependencies without explicit recurrence. The Anomaly Transformer [5] uses self-attention to learn pairwise associations between time points: it models both “prior” (local Gaussian-weighted) and “series” (learned self-attention) associations, and anomalies are detected when these expected associations are disrupted. Related models like TranAD [4] demonstrate that transformer-based models can differentiate anomalies by their deviant attention patterns and reconstruction discrepancies.
@@
xLSTM is rapidly gaining traction in different tasks. The most natural application as a direct successor of the LSTM architecture is in time series prediction tasks. The work in [14] proposed an advanced stock price prediction model based on xLSTM, with the aim of enhancing predictive accuracy in both short and long term periods. Experimental results with multiple stocks showed that xLSTM consistently outperforms LSTM across all stocks and time horizons, and this gap widens as the prediction period extends.
The success of xLSTM in time series prediction led researchers to make efforts towards its adaptation towards other tasks, data types, and domains. An example is lossless compression based on neural networks. The work in [8] propose a novel lossless compressor with two compression stages: parallel expansion mapping, which maps redundant pieces in multi-source data into unused alphabet values, and compression through an xLSTM model with a Deep Spatial Gating Module (DSGM). Authors in [15] introduce Vision-LSTM (ViL), an adaptation of xLSTM building blocks to computer vision. The architecture provides a stack of xLSTM blocks in which odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. The proposed model achieved strong classification performance in transfer learning and segmentation tasks, showing promise for further deployment of xLSTM in computer vision architectures. Similarly, LLMs built on the xLSTM architecture have emerged as a viable alternative to Transformers, due to a linear relationship between sequence length and compute requirements, as well as the constant memory usage. The authors in [9] introduced xLSTM 7B, a 7-billion-parameter LLM with targeted optimizations for fast and efficient inference. The model was able to achieve performance on downstream tasks similar to popular models such as Llama- and Mamba-based LLMs with a significantly faster inference speed [9].
@@ II-C TSB-AD anomaly detection benchmark
In the original TSB-AD study [1], a comprehensive collection of 40 time-series anomaly detection algorithms is leveraged to compare statistical, neural network-based, and the latest foundation model-based methods. Extensive experiments conducted on the TSB-AD benchmark led to relevant insights. First, statistical methods generally exhibit strong and consistent performance. In contrast, neural network-based models often fall short of their reputed superiority, though they still perform competitively in detecting point anomalies and handling multivariate time series. Moreover, simple architectures like CNNs and LSTMs tend to outperform more sophisticated designs such as advanced Transformer variants.
Even though CNN [2] and OmniAnomaly [13] display the strongest performance on multivariate datasets, the analysis provided by the TSB-AD benchmark highlights the need for more expressive models in complex multivariate settings.
@@ III Method
Figure 1: Proposed xLSTMAD Encoder-Decoder model architecture. Input data is organized as multivariate time series $x_{1},x_{2},\dots,x_{W}$, where $W$ is the window size. The data fed to the Encoder is processed into hidden states $z_{1},z_{2},\dots,z_{W}$ through linear projection, GELU activation, and multiple residually stacked xLSTM blocks. The Decoder generates reconstructions xLSTMAD-Ror forecasts xLSTMAD-F, which are then leveraged in the computation of anomaly scores. Dashed lines denote residual connections between stacked blocks.
@@
The Extended Long Short-Term Memory (xLSTM) [6] model architecture presents a significant advancement over the classical LSTM by addressing three major limitations: i) inability to revise past storage decisions, ii) limited memory capacity due to scalar-valued cell states, and iii) lack of parallelizability stemming from sequential hidden-to-hidden state dependencies. Specifically, a standard LSTM updates its memory cell as:
@@ 2. Encoder: Residual xLSTM Blocks
[…]
@@
$\mathbf{H}_{\ell}^{(f)}=\text{FFN}_{\ell}(\mathbf{H}_{\ell}^{(s)})\$ (8)
@@ III-D2 Phase-Aware Shape Consistency Loss via SoftDTW
To address the limitations of pointwise losses such as MSE in time series modeling, we incorporate a Soft Dynamic Time Warping (SoftDTW) loss [16]. Unlike traditional losses that penalize exact time-aligned discrepancies, SoftDTW enables differentiable sequence alignment, allowing for temporal shifts and warping, which are common in multivariate time series domains such as sensor networks, human activity recognition, and ECG signal monitoring.
@@
$\begin{split}\text{softmin}(a_{1},a_{2},a_{3})&=-\gamma\cdot\log(\exp\left(-\frac{a_{1}}{\gamma}\right)+\\ &+\exp\left(-\frac{a_{2}}{\gamma}\right)+\exp\left(-\frac{a_{3}}{\gamma}\right))\end{split}$ (22)
@@ IV Experimental Setup
In our experiments, we leverage the recently proposed comprehensive TSB-AD benchmark, which tackles the multiple issues affecting previous anomaly detection evaluations, such as flawed datasets, biased evaluation measures, and inconsistent benchmarking practices [1]. The benchmark includes a diverse set of multivariate time series datasets, with different underlying domain characteristics. This variety allows us to provide a comprehensive evaluation across various real-world challenges.
IV-A Metrics
@@
Popular metrics in anomaly detection include threshold-based metrics (Precision, Recall, F1-Score) that allow the evaluation of the various aspects of the performance of the model. A commonly used extension of F1-Score is the PA-F1 metric, which utilizes point adjustment on predictions. F1 has also been extended to consider the sequential nature of time series. The event-based F1 that considers each anomaly segment as a distinct event, contributing exactly once as a true positive or a false negative. On the other hand, the R-based-F1 extends traditional evaluation metrics by incorporating components such as existence reward, overlap reward, and a cardinality factor. Moreover, Affiliation-F1 emphasizes the temporal closeness between predicted and actual anomaly segments, quantifying their alignment based on the temporal distance [1]. However, the values of these metrics vary significantly depending on the choice of anomaly threshold.
@@
To evaluate a model using its corresponding anomaly scores, the Area Under the Receiver Operating Characteristics curve (AUC-ROC) [17] is a much more desirable metric. It is defined as the area under the curve corresponding to TPR on the y-axis and FPR on the x-axis, as the threshold of the anomaly score varies. Another metric following a similar idea is the Area Under the Precision-Recall curve (PR-AUC [18]), which leverages Precision on the y-axis and Recall on the x-axis. Although these metrics overcome the limitation of threshold-based metrics, they only gather point-based anomalies.
@@
Two desiderata in time series are range detection, i.e., the anomaly detection algorithm should ideally detect every point in the anomalous sequence, and existence detection, i.e., the detection of a tiny segment of one anomalous sequence is preferred rather than missing out on the whole sequence. To address this problem, a promising direction was recently proposed in [19], where the AUC was extended for range-based anomaly detection.
@@
VUS-ROC and VUS-PR build a surface of TPR, FPR based on varying window lengths. The volume under the surface then represents a measure of AUC for various windows. Since many time series anomaly detection models are sensitive to different window-length setups, the adoption of VUS enables a more fair and robust estimation and comparison of model performance. Due to these powerful properties, we focus our experiments on the VUS family metrics while including the other classical metrics for completeness. A more formal definition of the VUS family of measures can be found in [19].
IV-B Reproducibility guidelines
We follow the exact guidelines provided by the authors of the TSB-AD benchmark to ensure reproducibility and fair comparison of future methods. Although we provide a brief description of the evaluation scheme in the following, a detailed description and statistics can be found in the original paper [1]. The hyperparameter tuning set consists of 15% of each dataset, while the rest of the data is used to train and evaluate anomaly detection methods. The evaluation data is divided into 180 time series across 17 datasets. Each evaluation time series is then divided into training and test data sets. The datasets, along with information on all the splits, are available in the original TSB-AD repository: https://github.com/TheDatumOrg/TSB-AD/.

### v2

@@ I Introduction
Anomaly detection in multivariate time series is a vital task across domains such as industrial monitoring, cybersecurity, human activity recognition, and space system diagnostics. While statistical techniques and classical machine learning models remain strong baselines, deep learning methods have emerged as powerful alternatives due to their ability to model complex temporal dependencies and high-dimensional patterns. Recent benchmarks like TSB-AD [liu2024elephant] reveal that even simple neural architectures (e.g., CNNs [munir2018deepant], LSTMs [ji2021novel]) can outperform more complex Transformer variants [tuli2022tranad, xuanomaly] in anomaly detection tasks—particularly in multivariate settings where expressive temporal modeling is key.
Most state-of-the-art deep learning approaches fall into two broad categories: reconstruction-based models (e.g., autoencoders, VAEs) that detect anomalies through high reconstruction error, and prediction-based models (e.g., LSTMs) that forecast future behavior and identify deviations. However, these models often struggle to balance long-term memory, fine-grained temporal resolution, and computational efficiency [liu2024elephant].
The recently proposed xLSTM architecture offers a promising research direction in addressing all these gaps [beck2025xlstm]. As a generalization of LSTM, it combines depth-wise convolutions, residual connections, and multi-scale gating mechanisms in a stackable, efficient structure. xLSTM has demonstrated success in time series forecasting [auer2025tirexzeroshotforecastinglong], lossless compression [ma2025multi], and even large-scale language modeling [beck2025xlstm7b], where its linear memory footprint and fast inference make it a viable Transformer alternative. Despite its growing popularity, no prior work has explored xLSTM for anomaly detection.
@@
To rigorously evaluate our approach, we leverage the TSB-AD-M benchmark [liu2024elephant], a comprehensive suite of 17 real-world datasets ranging from industrial control systems (e.g., SWaT, SMD) to spacecraft telemetry (e.g., SMAP, MSL) and human activity monitoring (e.g., OPP), designed to create extremely challenging conditions that are useful to assess the effectiveness of diverse models in diverse conditions. Our results show that the xLSTM encoder-decoder architecture achieves superior performance compared to classical and modern baselines.
@@ II-A Time series anomaly detection
Anomaly detection in time series encompasses a range of classical and modern methods. Traditional statistical approaches assume an underlying generative model of normal behavior and use statistical tests or thresholds to flag outliers. For example, one may fit an ARIMA or seasonal model and then apply hypothesis tests on residuals [aggarwal2017introduction]. Clustering and density-based methods form another classical category. These unsupervised approaches group similar time-series segments (e.g., via k-means or hierarchical clustering) and define a normal ,,profile” of data. Points that lie far from cluster centroids or form very small clusters are labeled anomalous [aggarwal2017introduction, breunig2000lof]. Likewise, distance-based techniques (e.g., k-NN, LOF) and density-based methods (e.g., DBSCAN) treat points in sparse regions of feature space as outliers  [aggarwal2017introduction, li2020copod].
In the deep learning era, reconstruction-based models have become popular for time-series anomaly detection. Autoencoders (AEs) are neural networks trained to reconstruct input sequences, and anomalies are identified by high reconstruction error. Sakurada and Yairi [sakurada2014anomaly] showed that even simple AEs can outperform linear PCA for detecting anomalies in multivariate time series. Modern variants (e.g., denoising AEs, convolutional AEs) follow the same principle. For instance, the USAD model [audibert2020usad] uses two adversarially-trained autoencoders: by amplifying reconstruction error, it more clearly separates anomalous from normal patterns in multivariate data. In general, these methods learn a low-dimensional embedding of the time series and score anomalies by how poorly the model can reconstruct a given sequence.
Generative deep models explicitly learn the probability distribution of normal time-series data. Variational autoencoders (VAEs) and GANs are common in this class. For example, OmniAnomaly [su2019robust] uses a recurrent VAE with planar normalizing flows to model multivariate time series: it learns a compact latent representation of normal behavior and flags anomalies when the reconstruction probability is low. GAN-based schemes similarly train a generative model on normal sequences and use the generator or discriminator to score anomalies (e.g., a low discriminator confidence or high reconstruction loss indicates an anomaly).
Recently, transformer and attention-based architectures have been applied to anomaly detection. Transformers process the entire sequence via self-attention, capturing long-range dependencies without explicit recurrence. The Anomaly Transformer [xuanomaly] uses self-attention to learn pairwise associations between time points: it models both “prior” (local Gaussian-weighted) and “series” (learned self-attention) associations, and anomalies are detected when these expected associations are disrupted. Related models like TranAD [tuli2022tranad] demonstrate that transformer-based models can differentiate anomalies by their deviant attention patterns and reconstruction discrepancies.
@@
xLSTM is rapidly gaining traction in different tasks. The most natural application as a direct successor of the LSTM architecture is in time series prediction tasks. The work in [fan2024advanced] proposed an advanced stock price prediction model based on xLSTM, with the aim of enhancing predictive accuracy in both short and long term periods. Experimental results with multiple stocks showed that xLSTM consistently outperforms LSTM across all stocks and time horizons, and this gap widens as the prediction period extends.
The success of xLSTM in time series prediction led researchers to make efforts towards its adaptation towards other tasks, data types, and domains. An example is lossless compression based on neural networks. The work in [ma2025multi] propose a novel lossless compressor with two compression stages: parallel expansion mapping, which maps redundant pieces in multi-source data into unused alphabet values, and compression through an xLSTM model with a Deep Spatial Gating Module (DSGM). Authors in [alkin2024vision] introduce Vision-LSTM (ViL), an adaptation of xLSTM building blocks to computer vision. The architecture provides a stack of xLSTM blocks in which odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. The proposed model achieved strong classification performance in transfer learning and segmentation tasks, showing promise for further deployment of xLSTM in computer vision architectures. Similarly, LLMs built on the xLSTM architecture have emerged as a viable alternative to Transformers, due to a linear relationship between sequence length and compute requirements, as well as the constant memory usage. The authors in [beck2025xlstm7b] introduced xLSTM 7B, a 7-billion-parameter LLM with targeted optimizations for fast and efficient inference. The model was able to achieve performance on downstream tasks similar to popular models such as Llama- and Mamba-based LLMs with a significantly faster inference speed [beck2025xlstm7b].
@@ II-C TSB-AD anomaly detection benchmark
In the original TSB-AD study [liu2024elephant], a comprehensive collection of 40 time-series anomaly detection algorithms is leveraged to compare statistical, neural network-based, and the latest foundation model-based methods. Extensive experiments conducted on the TSB-AD benchmark led to relevant insights. First, statistical methods generally exhibit strong and consistent performance. In contrast, neural network-based models often fall short of their reputed superiority, though they still perform competitively in detecting point anomalies and handling multivariate time series. Moreover, simple architectures like CNNs and LSTMs tend to outperform more sophisticated designs such as advanced Transformer variants.
Even though CNN [munir2018deepant] and OmniAnomaly [su2019robust] display the strongest performance on multivariate datasets, the analysis provided by the TSB-AD benchmark highlights the need for more expressive models in complex multivariate settings.
@@ III Method
Figure 1: Proposed xLSTMAD Encoder-Decoder model architecture. Input data is organized as multivariate time series $x_{1},x_{2},\dots,x_{W}$, where $W$ is the window size. The data fed to the Encoder is processed into hidden states $z_{1},z_{2},\dots,z_{W}$ through linear projection, GELU activation, and multiple residually stacked xLSTM blocks consisting of mLSTM and optional sLSTM layers. The Decoder generates reconstructions xLSTMAD-Ror forecasts xLSTMAD-F, which are then leveraged in the computation of anomaly scores. Dashed lines denote residual connections between stacked blocks.
@@
The Extended Long Short-Term Memory (xLSTM) [beck2025xlstm] model architecture presents a significant advancement over the classical LSTM by addressing three major limitations: i) inability to revise past storage decisions, ii) limited memory capacity due to scalar-valued cell states, and iii) lack of parallelizability stemming from sequential hidden-to-hidden state dependencies. Specifically, a standard LSTM updates its memory cell as:
@@ 2. Encoder: Residual xLSTM Blocks
where $\text{Conv1D}_{8}$ corresponds to one dimensional convolutional layer with kelner size = 8.
@@
$\mathbf{H}_{\ell}^{(f)}=\text{FFN}_{\ell}(\mathbf{H}_{\ell}^{(s)})\\$ (8)
@@ III-D2 Phase-Aware Shape Consistency Loss via SoftDTW
To address the limitations of pointwise losses such as MSE in time series modeling, we incorporate a Soft Dynamic Time Warping (SoftDTW) loss [cuturi2017soft]. Unlike traditional losses that penalize exact time-aligned discrepancies, SoftDTW enables differentiable sequence alignment, allowing for temporal shifts and warping, which are common in multivariate time series domains such as sensor networks, human activity recognition, and ECG signal monitoring.
@@
[…]
@@ IV Experimental Setup
In our experiments, we leverage the recently proposed comprehensive TSB-AD benchmark, which tackles the multiple issues affecting previous anomaly detection evaluations, such as flawed datasets, biased evaluation measures, and inconsistent benchmarking practices [liu2024elephant]. The benchmark includes a diverse set of multivariate time series datasets, with different underlying domain characteristics. This variety allows us to provide a comprehensive evaluation across various real-world challenges. In the following, we describe each dataset in more detail.
IV-A Datasets
The TSB-AD time series anomaly detection benchmark contains an extensive collection of 13 univariate and 20 multivariate public time-series anomaly detection datasets. In this study, we focus on the multivariate datasets:
  • GHL [filonov2016multivariate]: Includes operational data from three reservoirs, capturing temperature and water level metrics. Anomalies are characterized by shifts in maximum temperature or deviations in pump frequency.
  • Daphnet [bachlin2009wearable]: Comprises annotated accelerometer data from three sensors placed on Parkinson’s disease patients. The dataset focuses on detecting episodes of freezing of gait (FoG) during walking tasks.
  • Exathlon [jacob2020exathlon]: Derived from real-world traces collected over 2.5 months from a Spark cluster. Each anomaly instance includes ground truth labels for both the root cause interval and the effect interval.
  • Genesis [von2018anomaly]: A dataset from a portable pick-and-place demonstrator system that utilizes an air tank to operate gripping and storage components.
  • Opportunity (OPP) [roggen2010collecting]: Consist of human activity data. It includes motion sensor data recorded as subjects performed typical daily activities.
  • Server Machine Dataset (SMD) [su2019robust]: The dataset captures server metrics for anomaly detection, collected over a five-week period from 28 machines.
  • SWaT [mathur2016swat]: A secure water treatment dataset with readings from 51 sensors and actuators. Anomalies in the dataset correspond to attack scenarios.
  • PSM [abdulaal2021practical]: Contains system-level metrics collected from various application server nodes at eBay.
  • SMAP [hundman2018detecting]: Composed of telemetry data from NASA’s Soil Moisture Active Passive (SMAP) satellite.
  • MSL [hundman2018detecting]: Collected from the Curiosity Rover as part of Mars Science Laboratory (MSL) operations. The dataset includes telemetry data with annotated anomalies.
  • CreditCard [sharafaldin2018toward]: An intrusion detection dataset containing labeled network flows, including full packet payloads, traffic profiles, and flow labels.
  • GECCO [moritz2018gecco]: A water quality dataset focused on the monitoring of drinking water safety.
  • MIT-BIH Arrhythmia Database (MITDB) [goldberger2000physiobank]: Contains 48 half-hour two-channel ECG recordings from 47 individuals.
  • Supraventricular Database (SVDB) [greenwald1990improved]: Provides 78 half-hour ECG recordings that complement the MITDB by emphasizing supraventricular arrhythmias.
  • Long-Term Database (LTDB) [goldberger2000physiobank]: A collection of seven extended-duration ECG recordings (14–22 hours each), featuring manually verified beat annotations.
  • CATSv2 [CATSV2]: The second version of the Controlled Anomalies Time Series dataset, consisting of command sequences, external stimuli, and telemetry from a simulated complex system.
  • TAO [TAO]: It contains 575,648 records with 3 attributes that are collected from the Tropical Atmosphere Ocean project.
IV-B Metrics
@@
Popular metrics in anomaly detection include threshold-based metrics (Precision, Recall, F1-Score) that allow the evaluation of the various aspects of the performance of the model. A commonly used extension of F1-Score is the PA-F1 metric, which utilizes point adjustment on predictions. F1 has also been extended to consider the sequential nature of time series. The event-based F1 that considers each anomaly segment as a distinct event, contributing exactly once as a true positive or a false negative. On the other hand, the R-based-F1 extends traditional evaluation metrics by incorporating components such as existence reward, overlap reward, and a cardinality factor. Moreover, Affiliation-F1 emphasizes the temporal closeness between predicted and actual anomaly segments, quantifying their alignment based on the temporal distance [liu2024elephant]. However, the values of these metrics vary significantly depending on the choice of anomaly threshold.
@@
To evaluate a model using its corresponding anomaly scores, the Area Under the Receiver Operating Characteristics curve (AUC-ROC) [fawcett2006introduction] is a much more desirable metric. It is defined as the area under the curve corresponding to TPR on the y-axis and FPR on the x-axis, as the threshold of the anomaly score varies. Another metric following a similar idea is the Area Under the Precision-Recall curve (PR-AUC [davis2006relationship]), which leverages Precision on the y-axis and Recall on the x-axis. Although these metrics overcome the limitation of threshold-based metrics, they only gather point-based anomalies.
@@
Two desiderata in time series are range detection, i.e., the anomaly detection algorithm should ideally detect every point in the anomalous sequence, and existence detection, i.e., the detection of a tiny segment of one anomalous sequence is preferred rather than missing out on the whole sequence. To address this problem, a promising direction was recently proposed in [paparrizos2022volume], where the AUC was extended for range-based anomaly detection.
@@
VUS-ROC and VUS-PR build a surface of TPR, FPR based on varying window lengths. The volume under the surface then represents a measure of AUC for various windows. Since many time series anomaly detection models are sensitive to different window-length setups, the adoption of VUS enables a more fair and robust estimation and comparison of model performance. Due to these powerful properties, we focus our experiments on the VUS family metrics while including the other classical metrics for completeness. A more formal definition of the VUS family of measures can be found in [paparrizos2022volume].
IV-C Reproducibility guidelines
We follow the exact guidelines provided by the authors of the TSB-AD benchmark to ensure reproducibility and fair comparison of future methods. Although we provide a brief description of the evaluation scheme in the following, a detailed description and statistics can be found in the original paper [liu2024elephant]. The hyperparameter tuning set consists of 15% of each dataset, while the rest of the data is used to train and evaluate anomaly detection methods. The evaluation data is divided into 180 time series across 17 datasets. Each evaluation time series is then divided into training and test data sets. The datasets, along with information on all the splits, are available in the original TSB-AD repository: https://github.com/TheDatumOrg/TSB-AD/.